OpenStudy (anonymous):

What is the product written in scientific notation? 3*10^20/4*10^5

Vocaloid (vocaloid):

well, the way we usually handle these problems is to separate the exponential part from the non-exponential part start with 3/4 = ?

OpenStudy (anonymous):

0.75

Vocaloid (vocaloid):

ok, now let's work on the other part (10^20)/(10^5) = ?

OpenStudy (anonymous):

is it 16

Vocaloid (vocaloid):

um, not quite... when we have the same base, and we're doing division, we subtract the exponents 10^20/10^5 = 10^(20-5) = ?

OpenStudy (anonymous):

i dont know ; :(

Vocaloid (vocaloid):

20-5 = ?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

is it 7.5*10^15